A store is having a sale on almonds and jelly beans. For 8 pounds of almonds and 2 pounds of jelly beans, the total cost is $12. For 3 pounds of almonds and 8 pounds of jelly beans, the total cost is $31. Find the cost for each pound of almonds and each pound of jelly beans.
Answer by ankor@dixie-net.com(22740)

For 8 pounds of almonds and 2 pounds of jelly beans, the total cost is $12.
For 3 pounds of almonds and 8 pounds of jelly beans, the total cost is $31.
Find the cost for each pound of almonds and each pound of jelly bean
:\
\
let a = price of 1 lb of almonds
let j = price of 1 lb of jb
:
Write an equation for each statement
8a + 2j = 12
3a + 8j = 31
Multiply the 1st equation by 4,subtract the 2nd equation
32a + 8j = 48
3a + 8j = 31
---------------subtraction eliminates j, find a
29a + 0 = 17
a = 17/29
a = $.59 a pound for almonds
Find j
3(.59) + 8j = 31
1.77 + 8j = 31
8j = 31 - 1.77
8j = 29.23
j = 29.23/8
j = $3.65 a pound for jb
:
Check this in the first original equation
8(.59) + 2(3.65) = 12
4.72 + 7.30 = 12.02 ~ 12, for some reason the solutions were not integer cents
Maybe a typo somewhere in this problem